(t) = A, cos(500+B) Find the complex power S, = +by received by the source Find the complex power S, 4+by received by the res please also find a4(W) and b4(VAR) (got cut off in the pic), thanks!
Answer #1

VS - Al cos ( 500t+ B ) VS By 6 (45° 14 us= 6Cos( 500t+U5) = re B4 L=8mn Wh = 600 x 8 x 10 = un Input Impedance zin= 4+ (ca/lCurrent througth Denme Тc x 14 60.448 9 (26.53 ) o som vi 12 SR 5047444133 0.094)_9157 A complex Pocer Received By Resitter